OpenStudy (anonymous):

Verify/Establish: [(1+sec theta)/sec theta)]=[sin^2 theta/(1-cos theta)] Would it be better to begin from the LHS or the RHS?

OpenStudy (goformit100):

RHS and LHS will work. You procceed from RHS to LHS

OpenStudy (anonymous):

Okay, would I go about it be putting sec theta in it's form of 1/cos theta?

OpenStudy (goformit100):

LHS :- [(1+sec theta)/sec theta)] = 1/ sec theta + sec theta/sec theta = (cos theta + 1) = ( 1 + cos theta )

OpenStudy (goformit100):

[sin^2 theta/(1-cos theta)] = [(1 - cos^2 theta)/(1-cos theta)] = [(1 + cos theta)(1 - cos theta)/(1-cos theta)] = (1 + cos theta) ....................= LHS .......................proved

OpenStudy (goformit100):

Okay ?

OpenStudy (anonymous):

Okay, I'm just looking over your process. Sorry it's taking me such a long time.

OpenStudy (anonymous):

Thank you so much for showing me.

OpenStudy (goformit100):

No worries Ma'am. you can take your own time.

OpenStudy (goformit100):

Welcome it's my pleasure.